Circular No. 6817

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998E IN NGC 5161 Simon Woodings, Ralph Martin, and Andrew Williams report their

discovery of an apparent supernova (R ∼ 16.5) on Jan. 29.7 UT with the Perth-Lowell 0.61-m reflector in the course the Perth Astronomy Research Group’s automated supernova search at Perth Observatory. SN 1998E is located 3′′ east and 8′′ south of the center of NGC 5161 (galaxy center at α = 13^{h}29^{m}14^{s}.1, δ = −33^{o}10′30′′, equinox 2000.0). The last previous image of the field was taken on Jan. 9, with no object visible at this position to mag ∼ 19.

N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports that M.-T. Ruiz, University of Chile, obtained spectrophotometry (range 370–700 nm; resolution 1200) of SN 1998E with the CTIO 4-m Blanco telescope on Jan. 31.35 UT. The reduced spectrophotometry shows a strong blue continuum with prominent emission lines, consistent with a type-IIn supernova classification. The most prominent emission lines are Hα, Hβ, Hγ, and Hδ, which have sharp cores (< 300 km/s) superimposed on broad wings that extend out to 5000 km/s from the rest wavelength. The line cores are slightly blueshifted with respect to the H I velocity of the galaxy by ∼ 50 km/s. Weaker emission lines of [Fe II] from multiplets 42, 46, 48, 49, and 74 are also present in the region 500–650 nm. There is no evidence for He I at 587.6 nm, which is commonly seen in early-stage type-IIn supernovae. The only absorption lines present are from Ca II H+K.

COMET C/1997 J2 (MEUNIER-DUPOUY) Visual m_{1} estimates: Jan. 1.71 UT, 10.6 (M. Lehký, Hradec Králové,

Czech Rep., 25×100 binoculars); 17.72, 11.1 (V. Znojil, Brno, Czech Rep., 25×100 binoculars); 23.74, 10.4 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector); 29.73, 10.9 (M. Plsek, Lelekovice, Czech Rep., 0.35-m reflector).

COMET 69P/TAYLOR Visual m_{1} estimates: Jan. 17.77 UT, 11.9 (K. Hornoch, Lelekovice,

Czech Republic, 0.35-m reflector); 22.05, 12.4 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); 25.83, 12.3 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 28.64, 12.3 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 29.83, 12.3 (Hornoch).

1998 January 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green